OpenStudy (anonymous):

re enter question need find equations for lines parallel to the given line but passing through the given point y=-4x-5;(0,-1)

OpenStudy (anonymous):

given y=-4x-5 here the slope = m = -4 when two lines are parallel their slopes are equal therefore the slope for the required line is also -4 now you have slope = m = -4 and point = (0,-1) use point slope form to get the equation y-y1 = m(x-x1) x1 = 0 , y1 = -1 and m = -4 put all these values in this equation and get the required equation